OpenStudy (anonymous):

Will give a medal to most helpful!!!!!! Solve the system of equations. x-4y=5 x=3y-2 (_,_)

OpenStudy (anonymous):

First you pick the second equation cause the x is already isolated and place it in first equation in the place of x

OpenStudy (anonymous):

ok.......

OpenStudy (anonymous):

Then it goes like this: 3y-2-4y=5 Solving this: -7y=7 Y=-1

OpenStudy (anonymous):

2 is the y intercet, right?

OpenStudy (anonymous):

Now that you found y you place it in the second equation a and find x X=3.-1-2 X=6

OpenStudy (anonymous):

(6,-1)

OpenStudy (anonymous):

ok, so it would be (6,-1)

OpenStudy (anonymous):

thank you so much :D

OpenStudy (anonymous):

@DonutChip ok, how would you do it then?

OpenStudy (anonymous):

I'm pretty sure that's right, but please explain how you would do it and why my resolution is wrong

OpenStudy (anonymous):

nevermind lol i did my math wrong .-.

OpenStudy (anonymous):

ok lol, I am going to close this question.Thank you do much @Mrs.Allison

OpenStudy (anonymous):

That's ok, it is really important to expresse your opinion and point out whenever you feel something is wrong, I'm a teacher and I really liked your attitude. Good job @DonutChip

OpenStudy (anonymous):

You are welcome

OpenStudy (mrnood):

@Mrs.Allison your solution IS wrong 3y-2-4y=5 (is correct substitution) solving gives -y= 7 y=-7 and substituting back in second equation: x= 3(-7)-2 = -23 CHECK X BACK IN THE FIRST EQUATION: -23 - 4(-7)= 5 = correct therefore (-23,-7) is the solution